# Efficient Radar Modulation Recognition via a Noise-Aware Ensemble Neural Network

### v2

@@ IV-A Training Recognition Model
A training dataset was built to train the recognition model, generating 5,000 simulated signals for each of the 12 modulation schemes. Table IV lists the parameters used in the signal generation, where $f_{s}$, $f_{min}$, $L_{hs}$, $f_{hop}$, $L_{B}$, and $N_{sc}$ represent sampling frequency, fundamental frequency of the Costas code, hopping sequence length of the Costas code, frequency spacing of the Costas code, length of the Barker code, and the number of samples per carrier frequency cycle, respectively. Additionally, $U(\cdot,\cdot)$ denotes a uniform distribution.
The proposed NAEL framework consists of three neural networks, each trained individually whereas keeping the weights of the other networks constant during training. For the performance analysis, all classification models used in this study, including the NAEL-based model, were trained using the Adam optimizer [37]. Training was conducted for 100 epochs, resulting in 187,500 weight updates. The input data fed into all neural networks were normalized to have a mean of 0 and a variance of 1.
Figure 9: Confusion matrix of the proposed recognition model using the test dataset collected in the (a) first scenario, (b) second scenario (low transmit power), and (c) third scenario (very low transmit power).
@@ IV-B Test Dataset Generation
The test dataset for performance analysis was constructed by collecting signals in a real-world environment, as shown in Fig. 8. Radar signals with 12 modulation schemes were generated using MATLAB. The generated digital signals were converted to analog signals using the USRP N210 software-defined radio platform[38]. The converted analog signals were transmitted via an antenna connected to a transmitting USRP N210 and received by an antenna connected to a distant receiving USRP N210. Vertical antennas were used for both transmission and reception, positioned 13.5 m apart. The radar signals used in this experiment had a center frequency of 1,902.5 MHz. The collected signals were digitized at a sampling rate of 10 MHz using a USRP N210 and imported into a computer for modulation recognition.
The signals were recorded for three different scenarios by varying the transmission power to analyze performance under various SNR conditions. In each scenario, 100 signals were collected for each of the 12 modulation schemes. One scenario with relatively high signal power and two scenarios with low signal power were used for the test dataset generation. The estimated SNRs for the first, second, and third scenarios were approximately –4, –15, and –17 dB, respectively. Notably, in the experiments using the USRP, the exact SNR for each scenario could not be calculated. However, an approximate SNR can be estimated by assuming that the power of the collected signal $y(n)$ is the sum of the power of the signal $x(n)$ and that of the signal collected in an environment without transmitted signals (i.e., received data with only noise $w(n)$ presence).
Figure 10: Performance versus computational cost of various models using the test dataset collected in the (a) first scenario, (b) second scenario (low transmit power), (c) third scenario (very low transmit power).
@@ IV-C Analysis Results on Experimental Data
In the first scenario (as seen in Fig. 9(a)), where the signal power is relatively high, the proposed recognition model exhibits low confusion, resulting in high classification performance. However, as shown in Fig. 9(b), many misclassifications occur as the SNR decreases. In particular, distinguishing between Frank and P3, as well as P1 and P4 is difficult, as analyzed in [15]. When the signal power was extremely low, the classification performance for all modulation schemes decreased, as shown in Fig. 9(c).
@@
In the high SNR scenario shown in Fig. 10(a), the proposed model exhibited the highest recognition performance and the lowest computational cost compared to the other models. This is because the NAN determined that most of the input data were not significantly affected by noise. Consequently, recognition could be performed using only the PRN at a relatively low computational cost, without activating the ARN. In fact, the NAN of our proposed model activated the ARN 20 times out of 1,200 predictions in the first scenario.
@@
In a low SNR environment scenario, as shown in Fig. 10(b), the PCCs for all classification models are around 50%. Among them, our proposed model exhibits the highest classification performance compared to other models. The computational cost of our proposed model was also the lowest among those achieving high recognition accuracy within the comparison group. In the third scenario, with extremely low signal power as shown in Fig. 10(c), the PCCs of the classification models are around 35%. Here, the proposed model shows superior performance with a PCC approximately 2.5% higher than the best-performing comparison model. This is because the NAN identifies significant noise-affected data in the input, utilizing both PRN and ARN adaptively for the final decision. The proposed model also incurs lower computational costs. Efficiency is achieved by effectively using only the PRN for classification in cases where data are likely unaffected by noise, thereby avoiding unnecessary activation of ARN.

### v3

@@ IV-A Training Recognition Model
A training dataset was built to train the recognition model, generating 5,000 simulated signals for each of the 12 modulation schemes. Table V lists the parameters used in the signal generation, where $f_{s}$, $f_{min}$, $L_{hs}$, $f_{hop}$, $L_{B}$, $N_{sc}$, $k$, and $\Delta F$ represent sampling frequency, fundamental frequency of the Costas code, hopping sequence length of the Costas code, frequency spacing of the Costas code, length of the Barker code, the number of samples per carrier frequency cycle, the number of segments, and modulation bandwidth, respectively. Additionally, $U(\cdot)$ denotes a uniform distribution.
The proposed NAEL framework consists of three neural networks, each trained individually whereas keeping the weights of the other networks constant during training. For the performance analysis, all classification models used in this study, including the NAEL-based model, were trained using the Adam optimizer [40]. The training was conducted for 100 epochs, resulting in 187,500 weight updates. The input data fed into all neural networks were normalized to have a mean of 0 and a variance of 1. All experiments were conducted on a workstation equipped with an AMD Ryzen 9 5950X CPU and an NVIDIA GeForce RTX 3080 Ti GPU, using the PyTorch framework.
Figure 9: Illustration of measurement environment for test dataset generation under realistic conditions.
Figure 10: Confusion matrix of the proposed recognition model using the test dataset collected in the (a) first scenario (high-SNR case), (b) second scenario (low-SNR case), and (c) third scenario (very low-SNR case).
@@ IV-B Test Dataset Generation
The test dataset for performance analysis was constructed by collecting signals in a real-world environment, as shown in Fig. 9. In this experiment, universal software radio peripheral (USRP) N210 software-defined radios (SDRs)[41] were employed. SDRs provide flexible and programmable platforms capable of transmitting and receiving various signal types, making them suitable for diverse wireless applications such as radar modulation recognition. Although practical SDR devices like the USRP N210 typically have constraints such as limited sampling rates and potential hardware-induced noise or distortion, their versatility and ease of use still make them well-suited for radar modulation recognition research.
In our experiment, radar signals with 12 modulation schemes were generated using MATLAB. The generated digital signals were converted to analog signals using the USRP N210 SDR platform. The converted analog signals were transmitted via an antenna connected to a transmitting USRP N210 and received by an antenna connected to a distant receiving USRP N210. Vertical antennas were used for both transmission and reception, positioned 13.5 m apart. The radar signals used in this experiment had a center frequency of 1,902.5 MHz. The collected signals were digitized at a sampling rate of 10 MHz using a USRP N210 and imported into a computer for modulation recognition.
These collected signals inherently reflect real-world conditions, incorporating multiple factors that deviate from an idealized additive white Gaussian noise (AWGN)-only environment. As illustrated in Fig. 9, multipath interference occurs due to the presence of obstacles, causing signal reflections that alter channel characteristics. Additionally, the use of USRP N210 hardware introduces hardware-induced distortions such as phase noise and frequency offset, further differentiating the experimental setting from a controlled simulation. By capturing these diverse real-world effects, the test dataset enables a comprehensive evaluation of the robustness of the proposed method.
The signals were recorded for three different scenarios by varying the transmission power to analyze performance under various SNR conditions. In each scenario, 100 signals were collected for each of the 12 modulation schemes. The test dataset comprised three scenarios: one with a high-SNR condition and two with low- and very low- SNR conditions. The estimated SNRs for the first, second, and third scenarios were approximately $-4$, $-15$, and $-17$ dB, respectively. Notably, in the experiments using the USRP, the exact SNR for each scenario could not be calculated. However, an approximate SNR can be estimated by assuming that the power of the collected signal $y(n)$ is the sum of the power of the signal $x(n)$ and that of the signal collected in an environment without transmitted signals (i.e., received data with only noise $w(n)$ presence).
@@ IV-C Analysis Results on Experimental Data
TABLE VI: PCC and FLOPs of Various Recognition Models Using the Test Dataset Collected in the First Scenario
      Recognition Model       PCC (%)       MFLOPs
      LPI-Net       93.8       69
      LWCNN       98.0       507
      MobileNetV1       98.8       701
      MobileNetV2       98.7       416
      ShuffleNetV1 2.0x       98.8       692
      ShuffleNetV2 2.0x       98.8       766
      MNasNet 1.3       98.6       702
      Vision Transformer       98.8       5,540
      NAEL       99.1       402
TABLE VII: PCC and FLOPs of Various Recognition Models Using the Test Dataset Collected in the Second Scenario
      Recognition Model       PCC (%)       MFLOPs
      LPI-Net       24.8       69
      LWCNN       43.4       507
      MobileNetV1       49.5       701
      MobileNetV2       51.1       416
      ShuffleNetV1 2.0x       51.6       692
      ShuffleNetV2 2.0x       52.1       766
      MNasNet 1.3       54.5       702
      Vision Transformer       51.4       5,540
      NAEL       55.4       606
TABLE VIII: PCC and FLOPs of Various Recognition Models Using the Test Dataset Collected in the Third Scenario
      Recognition Model       PCC (%)       MFLOPs
      LPI-Net       23.6       69
      LWCNN       30.2       507
      MobileNetV1       34.8       701
      MobileNetV2       33.4       416
      ShuffleNetV1 2.0x       35.5       692
      ShuffleNetV2 2.0x       33.9       766
      MNasNet 1.3       35.9       702
      Vision Transformer       33.8       5,540
      NAEL       38.3       660
In the first scenario (as seen in Fig. 10(a)), where the signal power is relatively high, the proposed recognition model exhibits low confusion, resulting in high classification performance. However, as shown in Fig. 10(b), many misclassifications occur as the SNR decreases due to the increasing influence of noise on the signal characteristics. In particular, distinguishing between Frank and P3, as well as P1 and P4, becomes significantly challenging. This is because these modulation schemes rely heavily on precise phase transitions for differentiation, and in low SNR conditions, phase information is highly susceptible to noise-induced distortions[15]. As a result, the model may incorrectly classify these modulation types due to the increased similarity in their observed signal patterns. Additionally, when the signal power was extremely low, the classification performance for all modulation schemes decreased, as shown in Fig. 10(c). This degradation occurs because the noise level highly exceeds the signal power, making it difficult for the model to extract meaningful features from the received data. Consequently, the overall classification accuracy decreases, with more modulation schemes being misclassified due to indistinguishable modulation characteristics in a high-noise environment.
@@
In the high SNR scenario shown in Table VI, the proposed model exhibited the highest recognition performance and the lowest computational cost compared to the other models. At high SNR levels, most signals are classified correctly using the PRN alone, which results in low computational cost while maintaining high PCC. Therefore, small variations in FLOPs occur due to occasional ARN activation when the model detects uncertain cases, leading to minor fluctuations in PCC. This is because the NAN determined that most of the input data were not significantly affected by noise. Consequently, recognition could be performed using only the PRN at a relatively low computational cost, without activating the ARN. In fact, the NAN of our proposed model activated the ARN 20 times out of 1,200 predictions in the first scenario, leading to an average computational load of approximately 402 MFLOPs and an average inference latency of 5.01 ms.
In a low SNR environment scenario, as shown in Table VII, the PCCs for all classification models are around 50 %. Among them, our proposed model exhibits the highest classification performance compared to other models. At lower SNR levels, the NAEL framework activates the ARN more frequently to compensate for increased noise level, which leads to an overall rise in FLOPs. In this scenario, ARN activation significantly increased, resulting in an average computational load of about 606 MFLOPs and an average inference time of 8.44 ms. The computational cost of our proposed model was also the lowest among those achieving high recognition accuracy within the comparison group.
Furthermore, our proposed method outperformed the Vision Transformer in recognition performance despite having lower computational complexity. This is because the proposed model effectively captures local feature relationships, whereas the Vision Transformer relies on a global self-attention mechanism that may not be optimal for our dataset. In modulation recognition tasks, the key features in time-frequency images are often spatially correlated, meaning that important information is locally connected rather than widely dispersed. Conventional CNNs, such as those used in our model, excel at learning these local patterns hierarchically, making them more effective in such cases. Conversely, Vision Transformers divide the input into patches and learn global dependencies, which is beneficial for large datasets with widely scattered features but less effective when key information is densely connected[43]. Considering the size of our dataset, the self-attention mechanism in Vision Transformers struggles to generalize effectively, resulting in suboptimal recognition performance compared to our approach.
@@
In the third scenario, with extremely low signal power as shown in Table VIII, the PCCs of the classification models are around 35 %. Here, the proposed model shows superior performance with a PCC approximately 2.4 % higher than the best-performing comparison model. This is because the NAN identifies significant noise-affected data in the input, utilizing both PRN and ARN adaptively for the final decision. However, in extremely low SNR environments, the activation frequency of ARN increases as noise impact on the input data becomes more significant. In this extremely low SNR scenario, the ARN activation rate was highest, triggered approximately 694 times (57.8 %) out of 1,200 predictions, increasing the average computational load to about 660 MFLOPs and the average inference time to 9.35 ms. While the model maintains a balance between accuracy and efficiency, this increased invocation may impact latency-sensitive applications or energy-constrained systems. Nevertheless, due to its adaptive structure, the proposed model still maintains lower overall computational costs compared to conventional models, as ARN is only activated when necessary. Efficiency is achieved by effectively using only the PRN for classification in cases where data are likely unaffected by noise, thereby avoiding unnecessary activation of ARN.
Figure 11: ARN call count distribution across different modulation schemes in the third scenario.
Additionally, we examined whether the noise awareness capability of NAEL is affected by the diversity of data. Fig. 11 illustrates the ARN call count distribution across different modulation schemes in the third scenario. The results show that while the call count varies, it does not exhibit a significant bias toward any specific modulation scheme. This suggests that while some variation exists, NAEL’s noise awareness capability remains relatively stable across different modulation schemes. Moreover, since our dataset encompasses various noise and interference environments, as presented in Fig. 9, the proposed method demonstrates robustness across diverse signal conditions, supporting its applicability in real-world scenarios.